A project requires an initial cash outflow of $16000. The investment is expected to produce net cash inflows of $2000 each year for eleven years. The internal rate of return on this investment to the nearest tenth of a percent is A. 5.7% B. 6.0% C. 5.9% D. 5.2%
Answer #1

Answer : A. 5.7%

=> Present value of cash inflow at 5% discount rate = cash inflow * PVIFA(5%,11) = 2000 * 8.306 = $16612

Present value of cash inflow at 6% discount rate = cash inflow * PVIFA(6%,11) = 2000 * 7.887 = $15774

Internal Rate of Return (IRR) = 5% + (16612 - 16000)/(16612 - 15774) * (6-5)% = 5% + (612/838) * 1% = 5% + 0.73% = 5.73%